# Bell nonlocality and entanglement in $\chi_{cJ}$ decays into baryon pair

### v1

@@ Abstract
We present a systematic analysis of Bell nonlocality and entanglement in $\chi_{cJ}$ decays into baryon pair ($B\bar{B}$) ($J=0,1,2$). From the baryon-antibaryon spin density matrix, we construct measurable Bell observables and concurrence, revealing a striking hierarchy of quantum correlations: $\chi_{c0}$ decays exhibit maximal violation and entanglement; $\chi_{c1}$ decays violate Bell inequalities for $\theta_{1}\in(0,\pi)$ with angle-modulated strength; we find that $B\bar{B}$ in $\chi_{c2}$ decays is in separable state, and no indication is found for the Bell inequality violation. We provide complete analytical results for $J=0,1$ and quantitative, uncertainty-aware estimations for $J=2$ using experimental inputs. The results indicates that the $\chi_{cJ}$ system as a novel platform for testing quantum entanglement in high-energy collisions.
@@ I Introduction
The $\chi_{cJ}$ states ($J=0,1,2$), with their distinct $J^{PC}$ quantum numbers ($0^{++},1^{++},2^{++}$), provide a rich laboratory in which the initial spin-parity structure is imprinted onto the spin correlations of the final-state baryon-antibaryon pairs. A central and open question is how these differing quantum numbers govern the violation of Bell inequalities and the degree of entanglement. In this work, we provide a comprehensive answer through a complete analysis of Bell nonlocality and spin entanglement in $\chi_{cJ}\to B\bar{B}$ decays.
Our analysis reveals a striking hierarchy of nonlocality across the $\chi_{cJ}$ states. We find that $\chi_{c0}$ decays maximally violate the Bell inequality. In contrast, $\chi_{c1}$ decays violate the inequality throughout the range $\theta_{1}\in(0,\pi)$, with the violation vanishing only at the exact forward and backward directions ($\theta_{1}=0,\pi$). The case of $\chi_{c2}$ is the most intricate, and the $B\bar{B}$ pair is in separable state; no indication is fournd for BI violation. These pronounced differences, directly rooted in the $J^{PC}$ quantum numbers of the parent $\chi_{cJ}$, yield clear and testable predictions for experiment.
The remainder of this paper is organized as follows. Section II introduces the theoretical framework for the joint spin density matrix. In Section III, we present the derived polarizations and spin correlations for the $B\bar{B}$ system in $\chi_{cJ}$ decays. Sections IV and V are devoted to the analysis of Bell nonlocality and quantum entanglement, respectively, where we provide analytical results and discuss their physical origins. Section VI presents a comprehensive numerical and phenomenological study: we first discuss the input parameters, then map out the landscape of Bell nonlocality and entanglement, and finally synthesize the physical implications. We conclude with a summary and outlook in Section VII.
@@ IV Bell nonlocality in $\chi_{cJ}\to B\bar{B}$ Decays
Bell’s theorem, through the derivation of the Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser:1969ny , provides a quantitative test to distinguish quantum mechanics from local hidden variable theories. In the Bell test with $\chi_{cJ}\to B\bar{B}$ decays, the experimentally measured spin correlation data are used to construct the correlation matrix $C$ in Eq. (1). The symmetric matrix $M=CC^{T}$ is then formed, and its eigenvalues $m_{1}\geq m_{2}\geq m_{3}$ are computed. According to the Horodecki condition Horodecki:1995nsk , the Bell inequality $I_{2}\leq 2$ is violated if and only if $m_{1}+m_{2}>1$. The maximum value of the CHSH observable for a given quantum state is given by
@@
$\mathbf{m}_{12}$ $=1+(\frac{(2-r_{1}^{2})\sin^{2}\theta_{1}}{2r_{1}^{2}+(2-r_{1}^{2})\sin^{2}\theta_{1}})^{2},~~~\mathrm{for~\chi_{c1}},$ (24)
@@
$m_{2}$ $=\frac{1}{2\mathcal{F}_{1}^{2}}(\mathcal{F}_{1}^{2}+\mathcal{F}_{2}^{2}+\mathcal{F}_{3}^{2}-4\mathcal{F}_{1}\mathcal{F}_{4}+2\mathcal{F}_{3}\mathcal{F}_{4}+5\mathcal{F}_{4}^{2}+\mathcal{F}_{2}^{2}\cos 2\Delta\Phi$
$-\sqrt{(\mathcal{F}_{1}+\mathcal{F}_{3}-\mathcal{F}_{4})^{2}((-\mathcal{F}_{1}+\mathcal{F}_{3}+3\mathcal{F}_{4})^{2}+2\mathcal{F}_{2}^{2}\cos(2\Delta\Phi)+2\mathcal{F}_{2}^{2})})$
$m_{3}$ $=\frac{1}{2\mathcal{F}_{1}^{2}}(\mathcal{F}_{1}^{2}+\mathcal{F}_{2}^{2}+\mathcal{F}_{3}^{2}-4\mathcal{F}_{1}\mathcal{F}_{4}+2\mathcal{F}_{3}\mathcal{F}_{4}+5\mathcal{F}_{4}^{2}+\mathcal{F}_{2}^{2}\cos 2\Delta\Phi$
$+\sqrt{(\mathcal{F}_{1}+\mathcal{F}_{3}-\mathcal{F}_{4})^{2}((-\mathcal{F}_{1}+\mathcal{F}_{3}+3\mathcal{F}_{4})^{2}+2\mathcal{F}_{2}^{2}\cos(2\Delta\Phi)+2\mathcal{F}_{2}^{2})}).$ (25)
@@ VI.1 Input parameters
$\chi_{c2}\to\Lambda\bar{\Lambda}$ $0.405\pm 0.009$ $x=1.74\pm 0.15$ $0.37\pm 0.16$ BESIII:2025gof
@@ VI.3 Profile of quantum entanglement
In this work, we have presented a comprehensive theoretical analysis of Bell nonlocality and quantum entanglement in the decays $\chi_{cJ}\to B\bar{B}$ ($J=0,1,2$). By deriving the joint spin density matrix for the baryon-antibaryon system within the helicity formalism, we computed the complete set of polarization and spin correlation parameters. From these, we obtained analytical formulas for the Bell measure $\mathbf{m}_{12}$ and the concurrence $\mathcal{C}[\rho]$, revealing a striking and hierarchical picture of quantum correlations across the $\chi_{cJ}$ family.
The $\chi_{c0}$ decay yields a $B\bar{B}$ pair in a pure, maximally entangled spin‑triplet state ($\mathcal{C}[\rho]=1$), which guarantees a maximal violation of the Bell inequality ($\mathbf{m}_{12}=2$). In contrast, the $\chi_{c1}$ decay produces a state with persistent, yet non‑maximal, entanglement and Bell nonlocality that depends on the angular distribution; violation is observed over a wide angular range and disappears only exactly in the forward and backward directions, where transverse spin correlations are quenched. Within the uncertainty of the estimated helicity values for $\chi_{c2}\to B\bar{B}$, we find that the $B\bar{B}$ pair is in a separable state, and no indication of Bell‑inequality violation is observed. Using the Horodecki criterion and the quantum concurrence, we show their dependence on the helicity ratio $x$ as presented in Tab.3.
@@
Looking forward, this analysis provides clear, measurable benchmarks for experiments at BESIII and future facilities such as the Super Tau-Charm Factory. In particular, the predicted angular distributions for $\chi_{c2}$ decays offer a direct means to extract the parameters $x$ and $\Delta\Phi$ more accurately. Precise measurement of these parameters will help infer quantum correlations via the mapping established in this work. Therefore, the $\chi_{cJ}$ system serves as a calibrated laboratory at the intersection of quantum information science and high-energy physics. Furthermore, studying quantum correlations in charmonium decays opens a new window into foundational quantum mechanics, enabling tests at energy scales and in regimes that are far removed from traditional optical experiments.

### v2

@@ Abstract
We present a systematic analysis of Bell nonlocality and entanglement in $\chi_{cJ}$($J=0,1,2$) decays into baryon pair($B\bar{B}$), with particular emphasis on their production via the process $e^{+}e^{-}\to\psi(2S)\to\gamma\chi_{cJ}$ at BESIII. From the baryon-antibaryon spin density matrix, we construct measurable Bell observables and concurrence, revealing a striking hierarchy of quantum correlations: $\chi_{c0}$ decays exhibit maximal violation and entanglement; $\chi_{c1}$ decays violate Bell inequalities for $\theta_{1}\in(0,\pi)$ with angle-modulated strength; we find that the $B\bar{B}$ pair in $\chi_{c2}$ decays is in a separable state, and no indication of Bell inequality violation is observed. We provide complete analytical results for $J=0,1$ and quantitative, uncertainty-aware estimations for $J=2$ based on experimental inputs from BESIII. These results establish the $\chi_{cJ}$ system produced via this radiative transition as a novel and promising platform for testing quantum entanglement and Bell nonlocality in high-energy collisions.
@@ I Introduction
The $\chi_{cJ}$ states ($J=0,1,2$), with their distinct $J^{PC}$ quantum numbers ($0^{++},1^{++},2^{++}$), provide a rich laboratory in which the initial spin-parity structure is imprinted onto the spin correlations of the final-state baryon-antibaryon pairs. In this work, we focus on $\chi_{cJ}$ mesons produced via the radiative transition $\psi(2S)\to\gamma\chi_{cJ}$ in $e^{+}e^{-}$ annihilation, which is the primary production mechanism accessible at BESIII with high statistics. A central and open question is how these differing quantum numbers govern the violation of Bell inequalities and the degree of entanglement. This work presents a comprehensive analysis of Bell nonlocality and spin entanglement in $\chi_{cJ}\to B\bar{B}$ decays, conducted under the well-defined production condition established here.
Our analysis reveals a striking hierarchy of nonlocality across the $\chi_{cJ}$ states. We find that $\chi_{c0}$ decays maximally violate the Bell inequality. In contrast, $\chi_{c1}$ decays violate the inequality throughout the range $\theta_{1}\in(0,\pi)$, with the violation vanishing only at the exact forward and backward directions ($\theta_{1}=0,\pi$). The case of $\chi_{c2}$ is the most intricate: the $B\bar{B}$ pair is found to be in a separable state, and no indication of Bell inequality violation is observed. These pronounced differences, directly rooted in the $J^{PC}$ quantum numbers of the parent $\chi_{cJ}$, yield clear and testable predictions for experiment.
The remainder of this paper is organized as follows. Section II introduces the theoretical framework for the joint spin density matrix. In Section III, we present the derived polarizations and spin correlations for the $B\bar{B}$ system in $\chi_{cJ}$ decays. Sections IV and V are devoted to the analysis of Bell nonlocality and quantum entanglement, respectively, where we provide analytical results and discuss their physical origins. Section VI presents a numerical analysis of the predicted quantum correlations using experimental inputs, followed by a detailed discussion of the experimental feasibility and future prospects for testing our predictions at BESIII and other facilities. We conclude with a summary and outlook in Section VII.
@@ IV Bell nonlocality in $\chi_{cJ}\to B\bar{B}$ Decays
Bell’s theorem, through the derivation of the Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser:1969ny , provides a quantitative criterion to assess whether the correlations exhibited by a quantum state are compatible with local hidden variable theories. A violation of the CHSH inequality ($\mathbf{m}_{12}>1$) signals that, under the premise of the quantum mechanics framework is accepted, the correlation predicted by the quantum state cannot be reproduced by any local hidden variable theory Horodecki:1995nsk .
However, the test of CHSH inequality at colliders remains an open question regarding its ability to refute local realism, primarily due to fundamental differences from optical Bell test setups. Collider experiments rely on quantum state tomography to reconstruct the spin density matrix from angular distributions of decay products, rather than allowing experimenters to freely and independently choose measurement settings. This absence of tunable optical inputs introduces an element of trust in the experimental platform, meaning the results constitute a witness of Bell non-locality rather than a true, device-independent Bell test. The effectiveness and significance of this loophole are different in collider experiments from those in optical experiments, and have been revisited in Refs Fabbrichesi:2024rec ; Fabbrichesi:2025aqp ; Shi:2019kjf . As reported recently in a BESIII measurement, the high significance of $\Lambda\bar{\Lambda}$ spin correlations has been observed, addressing the locality loophole by exploiting the weak decay distances of hyperons to achieve spacelike separation BESIII:2025vsr .
Nevertheless, showing that a quantum state violates a Bell inequality is still a significant physical result. It tells us that the spin correlations in a decay process cannot be explained by local realism. As a unique quantum feature, entanglement is not only widely applied in quantum science but also demonstrates novel value in high-energy particle physics: it can serve as a sensitive quantum probe in the search for evidence of CP violation Du:2024sly . In this spirit, we do not claim a definitive loophole-free test of the CHSH inequality. Instead, we use the helicity amplitudes measured by BESIII to see if the $\chi_{cJ}\to B\bar{B}$ system shows signs of Bell nonlocality. This gives us a concrete, data-driven foundation for planning future experiments.
To assess whether the $\chi_{cJ}\to B\bar{B}$ system exhibits Bell nonlocality, the experimentally measured spin correlation data are used to construct the correlation matrix $C$ in Eq. (1). The symmetric matrix $M=CC^{T}$ is then formed, and its eigenvalues $m_{1}\geq m_{2}\geq m_{3}$ are computed. According to the Horodecki condition Horodecki:1995nsk , the Bell inequality $I_{2}\leq 2$ is violated if and only if $m_{1}+m_{2}>1$. The maximum value of the CHSH observable for a given quantum state is given by
@@
$\mathbf{m}_{12}$ $=1+\left(\frac{(2-r_{1}^{2})\sin^{2}\theta_{1}}{2r_{1}^{2}+(2-r_{1}^{2})\sin^{2}\theta_{1}}\right)^{2},~~~\mathrm{for~\chi_{c1}},$ (24)
@@
$m_{2}$ $=\frac{1}{2\mathcal{F}_{1}^{2}}\left(\mathcal{F}_{1}^{2}+\mathcal{F}_{2}^{2}+\mathcal{F}_{3}^{2}-4\mathcal{F}_{1}\mathcal{F}_{4}+2\mathcal{F}_{3}\mathcal{F}_{4}+5\mathcal{F}_{4}^{2}+\mathcal{F}_{2}^{2}\cos 2\Delta\Phi\right$.
$\quad\left.-\sqrt{(\mathcal{F}_{1}+\mathcal{F}_{3}-\mathcal{F}_{4})^{2}\left((-\mathcal{F}_{1}+\mathcal{F}_{3}+3\mathcal{F}_{4})^{2}+2\mathcal{F}_{2}^{2}\cos(2\Delta\Phi)+2\mathcal{F}_{2}^{2}\right)}\right),$
$m_{3}$ $=\frac{1}{2\mathcal{F}_{1}^{2}}\left(\mathcal{F}_{1}^{2}+\mathcal{F}_{2}^{2}+\mathcal{F}_{3}^{2}-4\mathcal{F}_{1}\mathcal{F}_{4}+2\mathcal{F}_{3}\mathcal{F}_{4}+5\mathcal{F}_{4}^{2}+\mathcal{F}_{2}^{2}\cos 2\Delta\Phi\right$.
$\quad\left.+\sqrt{(\mathcal{F}_{1}+\mathcal{F}_{3}-\mathcal{F}_{4})^{2}\left((-\mathcal{F}_{1}+\mathcal{F}_{3}+3\mathcal{F}_{4})^{2}+2\mathcal{F}_{2}^{2}\cos(2\Delta\Phi)+2\mathcal{F}_{2}^{2}\right)}\right).$ (25)
@@ VI.1 Input parameters
$\chi_{c2}\to\Lambda\bar{\Lambda}$ $0.211\pm 0.112$ $x=1.74\pm 0.15$ $0.37\pm 0.16$ BESIII:2025gof
@@ VI.3 Profile of quantum entanglement
VI.4 Feasibility and Prospects
Based on the numerical results obtained from current experimental inputs, we now assess the feasibility of observing the predicted quantum correlations. This assessment takes into account the existing BESIII data sets and other facilities.
VI.4.1 Current Data and Future Precision
BESIII has accumulated the world’s largest sample of $\psi(2S)$ events. For $\psi(2S)\to\gamma\chi_{cJ},~\chi_{cJ}\to\Lambda\bar{\Lambda}$ decay, BESIII has recently published a high-statistics partial wave analysis using the full $\psi(2S)$ data sample of $2.712\times 10^{9}$ events BESIII:2025gof . Both the helicity amplitude ratios and the relative phase angles are measured with high precision, and systematic uncertainties dominate the error budget (e.g., $R_{\chi_{c2}}=0.575\pm 0.048\pm 0.018$, $\Delta\Phi_{\chi_{c2}}=-0.37\pm 0.15\pm 0.05$ rad). Consequently, the extracted $\mathbf{m}_{12}$ and concurrence for this channel are sufficiently precise to conclusively verify that the Horodecki condition is satisfied.
For $\chi_{cJ}\to p\bar{p}$ and $\chi_{cJ}\to\Xi\bar{\Xi}$ decays, the BESIII measurements are only based on partial data sets ($p\bar{p}$: $1.06\times 10^{8}$ $\psi(2S)$ events BESIII:2013nam , $\Xi^{-}\bar{\Xi}^{+}$, $\Xi^{0}\bar{\Xi}^{0}$: $4.481\times 10^{8}$ $\psi(2S)$ events BESIII:2022mfx ). The branching fractions of these decay modes are of the same order as that of $\chi_{cJ}\to\Lambda\bar{\Lambda}$ (typically $\mathcal{O}(10^{-4}\sim 10^{-5})$). Therefore, with the full $2.712\times 10^{9}$ $\psi(2S)$ data sample available at BESIII, the signal yields for $p\bar{p}$ and $\Xi\bar{\Xi}$ as well as $\Sigma\bar{\Sigma}$ final states are expected to be comparable to those of the $\Lambda\bar{\Lambda}$ channel. By scaling the systematic uncertainties from the $\Lambda\bar{\Lambda}$ measurement, we estimate that the precision of helicity parameters for these channels could be improved by a factor of 2233 once the full data set is analyzed. This would significantly reduce the error bands as shown in the figures of these decay modes, and enable to draw definitive conclusions on their Bell nonlocality and entanglement properties.
VI.4.2 Prospects at other facilities
Building on the above analysis, we next examine the feasibility of investigating $\chi_{cJ}\to B\bar{B}$ decays at other high-energy collider experiments.
LHCb and the Belle experiments offer valuable opportunities for heavy-flavor physics, but two challenges are confronted to probe $\chi_{cJ}\to B\bar{B}$ decays: low signal yield and the complexity of reconstructing the spin density matrix.
At LHCb, $\chi_{cJ}$ states are predominantly produced via $b$-hadron decays ($B\to\chi_{cJ}K$) LHCb:2013mbz or prompt production in $pp$ collisions LHCb:2013ofo , but the yield is less than a tenth of that at BESIII. At the Belle experiments, which operates at the $\Upsilon(4S)$ resonance Belle-II:2018jsg ; Belle-II:2010dht , the production rate via an initial state radiation (ISR) process, $e^{+}e^{-}\to\gamma_{ISR}\psi(2S)$, and $\psi(2S)\to\gamma\chi_{cJ}$ is also severely suppressed, with yields orders of magnitude below BESIII’s $2.7\times 10^{9}$ sample Belle:2015hcs . The both experiments face a fundamental challenge to reconstruct the $\chi_{cJ}$ spin density matrix. Compared to $e^{+}e^{-}$ annihilation at the BESIII, the $\chi_{cJ}$ mesons may be abundantly produced via other mechanisms at both facilities, but their spin density matrices are not directly accessible. One has to do a partial wave analysis to extract the helicity amplitudes of $\chi_{cJ}$ states by considering different production mechanisms. Nevertheless, if these matrices can be experimentally reconstructed, the same quantum tomography framework remains applicable.
The proposed Super $\tau$-Charm Factory (STCF) in China is designed to operate in the $\tau$-charm energy region with a peak luminosity of $0.5\times 10^{35}~\text{cm}^{-2}\text{s}^{-1}$ — two orders of magnitude higher than BEPCII Achasov:2023gey . With such a facility, the production rate of $\psi(2S)\to\gamma\chi_{cJ}$ events would be increased by a factor of $\sim 100$. With sufficiently high statistics, it would reduce statistical uncertainties to the sub-percent level, and the measurement precision would be dominated by systematic uncertainties, enabling percent-level tests of quantum correlation and measurements of helicity parameters for all ground baryonic pairs. If a polarized electron beam is used at the STCF in the future, the hyperons will acquire a longitudinal polarization. This will significantly enhance the spin analysis power for hyperons Salone ; Zeng , thereby providing more favorable conditions for studying the polarization correlation of baryon pairs and for testing Bell’s inequality with polarized beams Zhang:2026nwm . This would address the fundamental limitations discussed in Sec. IV. Therefore, the STCF represents the definitive future facility for the kind of quantum foundation studies advocated in this work.
@@ VII Summary and Outlook
In this work, we have presented a comprehensive theoretical analysis of Bell nonlocality and quantum entanglement in the decays $\chi_{cJ}\to B\bar{B}$ ($J=0,1,2$), specifically for $\chi_{cJ}$ mesons produced via the radiative transition $\psi(2S)\to\gamma\chi_{cJ}$ in $e^{+}e^{-}$ annihilation. By deriving the joint spin density matrix for the baryon-antibaryon system within the helicity formalism, we computed the complete set of polarization and spin correlation parameters. From these, we obtained analytical formulas for the Bell measure $\mathbf{m}_{12}$ and the concurrence $\mathcal{C}[\rho]$, revealing a striking and hierarchical picture of quantum correlations across the $\chi_{cJ}$ family under this well-defined production condition.
The $\chi_{c0}$ decay yields a $B\bar{B}$ pair in a pure, maximally entangled spin-triplet state ($\mathcal{C}[\rho]=1$), which guarantees a maximal violation of the Bell inequality ($\mathbf{m}_{12}=2$). In contrast, the $\chi_{c1}$ decay produces a state with persistent, yet non-maximal, entanglement and Bell nonlocality that depends on the angular distribution; violation is observed over a wide angular range and disappears only exactly in the forward and backward directions, where transverse spin correlations are quenched. Within the uncertainty of the estimated helicity parameters for $\chi_{c2}\to B\bar{B}$, we find that the $B\bar{B}$ pair is in a separable state, and no indication of Bell inequality violation is observed. Using the Horodecki criterion and the quantum concurrence, we show their dependence on the helicity ratio $x$ as presented in Table 3.
@@
The predictions in this work are specific to the production chain $e^{+}e^{-}\to\psi(2S)\to\gamma\chi_{cJ}$. They can be tested directly at BESIII using its full $\psi(2S)$ data sample. Looking forward, this analysis provides clear, measurable benchmarks for experiments at BESIII and future facilities such as the Super Tau-Charm Factory. In particular, the predicted angular distributions for $\chi_{c2}$ decays offer a direct means to extract the parameters $x$ and $\Delta\Phi$ more accurately. Precise measurement of these parameters will help infer quantum correlations via the mapping established in this work. A detailed assessment of the experimental feasibility, including the required event yields and the prospects at BESIII, LHCb, the Belle experiments, and future STCF, is also presented (see VI.4). These indicate that the $\chi_{cJ}$ system produced via this radiative transition serves as a calibrated laboratory at the intersection of quantum information science and high-energy physics. Furthermore, studying quantum correlations in charmonium decays opens a new window into foundational quantum mechanics, enabling tests at energy scales and in regimes that are far removed from traditional optical experiments.